OpenStudy (anonymous):

Can someone please explain the steps in order to solve this problem?!!!!? -2y+12=32+24y

OpenStudy (anonymous):

bring terms containing y on one side.. add... divide

OpenStudy (anonymous):

you bring the constants to one side and the variables to the other, so you add 2y to both sides and subtract 20 and you get -20 = 26y

OpenStudy (anonymous):

then you can divide and get the answer for y

OpenStudy (anonymous):

can you type in the steps. I am still confused

OpenStudy (anonymous):

My suggestion is to get the y terms to the left side of the equation and the constant terms on the right side of the equation using inverse operations. You can start by subtracting 12 from both sides.

OpenStudy (anonymous):

ok -2y=20+24y

OpenStudy (anonymous):

Now, you want all the y terms on the left side and also to keep the equation balanced. What do you think the next step should be? inverse operations: addition <-> subtraction multiplication <-> division

OpenStudy (anonymous):

divide the Y...but which side

OpenStudy (anonymous):

your not up to that step yet. You still need to separate the 24y from the 20.

OpenStudy (anonymous):

ok that is what I'm missing

OpenStudy (anonymous):

how to get rid of the 20?

OpenStudy (anonymous):

Not get rid of the 20, move the 24y to the other side by subtracting from both sides.

OpenStudy (anonymous):

-2y=20+24y -24 -24 ???????????????????

OpenStudy (anonymous):

Not -24 ---> -24y The whole term needs to be kept together.

OpenStudy (anonymous):

ok ok ok I think I'm getting it

OpenStudy (anonymous):

Great! What have you got so far?

OpenStudy (anonymous):

-26y=20

OpenStudy (anonymous):

Yes. What do you think the next step will be? Rembemer that you want to isolate the y.

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Is your answer expected to be in fraction or decimal form?

OpenStudy (anonymous):

-26y 20 ____ = ____ -26y -26y ??????????

OpenStudy (anonymous):

the answer is a fraction

OpenStudy (anonymous):

Almost. You want to separate the -26 from the y. So you will divide both sides by -26.

OpenStudy (uri):

-2y+12=32+24y Bring all y values on the left side and all numericals at one side Remember when a value changes it's side the sigh changes too So, -24y-2y=32-12 -26y=20 Now divide 20 with 26 So, y=20/-26 y=-0.769

OpenStudy (anonymous):

If your answer should be a fraction, make sure it is reduced to simplest form.

OpenStudy (anonymous):

it says the answer is - 11/13 ????????????????????????????????????

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

sorry it says -10/13

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

I don't understand how you get - 10/13

OpenStudy (anonymous):

Do you understand how \(y = -\frac{20}{26}\)?

OpenStudy (anonymous):

20 divided by 26?

OpenStudy (anonymous):

Divided by -26. Don't lose the negative.

OpenStudy (anonymous):

is a decimal .76

OpenStudy (anonymous):

Then that fraction needs to be reduced.

OpenStudy (anonymous):

\[-\frac{20}{26} \div \frac{2}{2} = -\frac{10}{13}\]

OpenStudy (anonymous):

@uri 26/2 = 13

OpenStudy (uri):

Yeah typo.

OpenStudy (uri):

The ansur is 0.76 like i explained before.

OpenStudy (anonymous):

@Algebra2013 How are you doing? Do you understand it?

OpenStudy (anonymous):

thank you gypsy!!!!!!!!!!!!!!!

OpenStudy (anonymous):

@uri Except that the answer needs to be a fraction.

OpenStudy (anonymous):

simplify 20/-26 you divide by 2.

OpenStudy (uri):

10/-13 or 0.76 same.

OpenStudy (anonymous):

Yes.

OpenStudy (uri):

He wants it in fraction so you have to write 10/-13...coool:)

OpenStudy (anonymous):

can we work out another problem??

OpenStudy (anonymous):

Sure.

OpenStudy (uri):

Sure...post in a new thread,Close this one.

OpenStudy (anonymous):

5x-12=13-4x

OpenStudy (anonymous):

Same steps. Do you know what to do first?

OpenStudy (uri):

@Algebra2013 Post this as a new question.Close this one

OpenStudy (anonymous):

add 12 to each side

OpenStudy (anonymous):

Yes.

OpenStudy (uri):

ONE QUESTION AT A TIME....if you don't want to get suspended.

OpenStudy (anonymous):

It is a continuation of the first question - Algebra simply wants confirmation that the process is understood.